|
Документ взят из кэша поисковой машины. Адрес
оригинального документа
: http://www.stsci.edu/smov/notes/0409smov.html
Дата изменения: Sat Apr 19 00:37:35 1997 Дата индексирования: Thu Feb 27 23:56:05 2014 Кодировка: Поисковые слова: http www.badastronomy.com bad tv foxapollo.html |
SMOV STATUS & PLANNING MEETING - 9 APR. 97
NOTE: The next meeting is April 11, 8:30 AM.
STATUS REPORTS
1. Observatory Status - All systems nominal.
Battery 6 reconditioning is proceeding nominally with the
discharge cycle complete.
2. WFPC2 Status - Proceeding nominally with Cycle 6 science and
remaining SMOV monitors.
3. NICMOS Status
- Engineering Status -
The NICMOS Dewar heaters suffered three more SAA-induced
thermal set point corruptions (two on April 7 and one on
April 8), The dewar heaters were commanded back off by the
STOCC per procedure.
The Detector temperature (mounting cups) is 61.0 degK, warming
at a rate of +0.02 degK/day.
The cold well (ice) temperature is 60.0 degK, warming at a rate
of +0.02 degK/day.
The VCS temperature is 98.5 degK. warming at a rate of +0.08
degK/day. On April 3 (093) the VCS started warming
up, but now appears to be stabilizing close to its current
temperature.
The cryogen venting had stabilized but now appears to be slightly
increasing again at a rate of ~ 156 grams/day. NICMOS has to date
lost an estimated total of 4% of its cryogen.
- Science Status - 7150 Focus monitor phase retrieval analysis
(Burrows, et al.) indicates no additional longitudinal shift
in focus positions. There are some small lateral shifts.
EROs 7116 (Arp220) and 7117 (M16) continue to be analyzed.
The Team is developing tools for data processing of individual
exposures and for updating pipeline inputs.
The 7034 Field Offset Mechanism (FOM) test was executed during the
night. It appeared to go nominally, though data analysis has not
yet begun.
4. STIS Status
- Engineering Status - Nominal.
- Science Status - Nominal. Analysis of the 7067 CCD target
acquisition test (3rd iteration) is ongoing. Like the second
iteration (as reported in the notes of 2 April), it was successful
in acquiring the target but the offset needed to go from
a filtered aperture to an unfiltered slit was not applied by the
on board algorithm.
Recents darks indicate that the warm-up resulting from last week's
Safemode "cleaned up" the hot pixels in the CCD.
The UV Monitor (7063) continues to indicate no loss in throughput,
although an apparent decline at the red end has been seen and is not
yet understood.
Vicki Balzano reported that, two days ago, it was discovered that
some PDB Symbol of Interest Files (SOIF) are incorrect, i.e, there
is a mismatch between memory locations specified by the SOIF and
the actual on-board memory locations. A Code 512
analysis has since narrowed the problem down to two flight software
changes made last year (while still under BASD control, prior
to 512 FSW control) for which the proper SOIF updates were not
made. This resulted in a misplacement of 25 onboard tables.
Planned corrective actions are for a reload of the memory in the
SMS111, adding an appropriate offset to each of the SOIF locations.
Then the SOIF files will be corrected in a later PDB release.
5. FGS status - No new activities.
6. FOC/COSTAR status- The FOC SMOV program is complete. The SMOV TTRB has
enabled the resumption of FOC science.
PLANNING REPORTS
1. Status of Upcoming SMSs
- SMS097 - Currenly executing.
- (Note: While composing these notes (afternoon of
9 April), I got word that HST has entered safemode.)
No details yet available, but the SMOV activities and
EROs scheduled over the next couple of days are obviously
impacted.)
- SMS104 - Final Mode products are expected to be available
tomorrow. Gaps for FSW configuration opportunities
related to MAMA SAA-avoidance management are on days 105
and 110.
- The CCD annealing proposal (7107) goes in this SMS.
- A PDB update involving the downlink transmission delay
parameter is in the hands of the PDB office, which
apparently is awaiting freeze board approval of the
change.
- Some edits have been made to manage the NICMOS dewar
heater set point through SAA-impacted orbits. The edits
will result in the refreshing/resetting of the heater
set point after every SAA passage. This will be
- SMS111 - Currently being built. There is a "slight
possibility" for the the MAMA turn-on (7105 revised) in
this SMS.
- SMS118 - Building has started. May include the bulk of
the MAMA turn on sequence (7105).
2. Replans
STIS
- MAMA HV Turn-on (7105) is currently planned for SMS111 or
SMS118. Downstream MAMA scheduling is being assessed.
NICMOS
- No significant replans have yet been identified.
OPERATIONS STATUS
1. OPUS - Nominal. No new problems. Fixes for previously reported
problems are coming in.
2. Archive - Nominal.